Radon has a half life of 56 seconds after what time will it’s activity be reduced to 90 percent
1
Expert's answer
2020-08-10T19:54:22-0400

The notion of half-life means that after 56 seconds the activity will decrease by 2.

"\dfrac{N(t)}{N_0} = 2^{-\lambda t}," here "0.5 = 2^{-t/56} \Rightarrow \lambda = 1/56."


"\dfrac{N(t)}{N_0} = 2^{- t/56}," so "\; 0.90 = 2^{-t/56} , \; 0.90 = e^{-(\ln 2)\cdot t/56} \Rightarrow t \approx 8.5\,\mathrm{s}."